OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = -3x + 12 and passes through the point (-1, 6). y = one thirdx + 7 y = -3x + 3 y = one thirdx + 3 y = -3x + 7

OpenStudy (anonymous):

@gorv

OpenStudy (gorv):

EQUATION OF PARALLEL LINE y=mx+c where m = slope of parallel line

OpenStudy (gorv):

parallel line have same slope so m=-3

OpenStudy (gorv):

y=-3x+c now it is passing through (-1,6)

OpenStudy (gorv):

plug in these point and find the value of c

OpenStudy (gorv):

6=-3*(-1)+c what u got ??

OpenStudy (anonymous):

-3?

OpenStudy (anonymous):

No wait. Positive 3.

OpenStudy (gorv):

6=3+c 6-3=c 3=c

OpenStudy (gorv):

yeah u r right

OpenStudy (gorv):

y=-3x+3 ie our equation

OpenStudy (anonymous):

Can you help with one more, PLEASE? :)

OpenStudy (gorv):

yeah :)